OpenStudy (anonymous):

The forces F 1, F 2, ..., F n acting at the same point P are said to be in equilibrium if the resultant force is zero, that is, if F 1 + F 2 + ... + F n = 0. F 1 = i + j, F 2 = i – j, F 3 = 9 i + 2 j Find the additional force required for the forces to be in equilibrium P.

OpenStudy (anonymous):

do the same as we did F1 =<1, 1> in this case, consider F1 as u = <u1,u2>=<1,1> F2= <1,-1> F2 as v = <v1,v2>= <1,-1> F3= <9,2> F3 as w =<w1,w2>=<9,2> so you need F4 = <x, y> such that x + u1+v1+w1 =0 or x = -u1-v1-w1 do the same with y combine them to get F4

OpenStudy (anonymous):

so F4 = <11,2>?

OpenStudy (anonymous):

<-11,-2> be careful with the signs

OpenStudy (anonymous):

??? x = u1+v1+w1 ? therefore 1+1+9 = 11 ?

OpenStudy (anonymous):

re read, I said x + u1+ v1 + w1 =0 or x = - u1 -v1-w1 = -1-1-9 =-11

OpenStudy (anonymous):

if x + u1 + v1 + w1 it will become 0?

OpenStudy (anonymous):

is it bcoz x has not value ?

OpenStudy (anonymous):

think of this situation, you apply a force to the right your girlfriend applies a force to the right your closed friend applies a force to the right I want to balance the force, so I must applies a force TO THE LEFT right? and the magnitude of my force must be = the sum of you guys, right?

OpenStudy (anonymous):

ahhhhh !! i got it ..

OpenStudy (anonymous):

good for you, hehehe

OpenStudy (anonymous):

so its -11,-2 ?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

oh come on !! youre my hero !! youre a lifesaver ^^

OpenStudy (anonymous):

ty